OpenStudy (anonymous):

Convert the temperature from degrees Fahrenheit to degrees Celsius, using the formula below. C = 5/9× (F – 32) 68°F = °C

OpenStudy (callisto):

Put F =68° intoC = 5/9× (F – 32) C = 5/8 (68-32) Can you do it from here?

OpenStudy (anonymous):

C=5/9 * 36

OpenStudy (callisto):

And simplify it?

OpenStudy (anonymous):

thats wehre i need help like do you need to divided 5/9 ?

OpenStudy (callisto):

C=5/9 * 36 = 5 x 36 /9 Can you simplify it now?

OpenStudy (anonymous):

yes i can wait

OpenStudy (anonymous):

180 and 45 ?

OpenStudy (callisto):

\[\frac{5}{9}\times 36 = \frac{5\times 36}{9} = \frac{5\times 4 \times 9}{9}\]What would you get from here?

OpenStudy (anonymous):

180/9

Parth (parthkohli):

That is equal to 20...

OpenStudy (anonymous):

and when you divided that then is going to be equal to 20

OpenStudy (callisto):